OpenStudy (ericamarie7):

WILL FAN & MEDAL!!!! Mrs. Cleary's class is selling candy bars to raise money for a field trip. The students in the class set a goal of how much money they would like to raise. The following formula describes this scenario: p=g/n where g=goal for money raised p=profit made from each candy bar sold n=number of candy bars sold The class wants to raise a total of $600.If they sell 600 candy bars, how much profit will they receive from each candy bar?

OpenStudy (cinnamonbun):

p=g/n p=600/600

OpenStudy (cinnamonbun):

wait wait